# Covariances for Free: Exploiting Mean Distributions for Federated Learning with Pre-Trained Models

### v1

@@ Abstract
Using pre-trained models has been found to reduce the effect of data heterogeneity and speed up federated learning algorithms. Recent works have investigated the use of first-order statistics and second-order statistics to aggregate local client data distributions at the server and achieve very high performance without any training. In this work we propose a training-free method based on an unbiased estimator of class covariance matrices. Our method, which only uses first-order statistics in the form of class means communicated by clients to the server, incurs only a fraction of the communication costs required by methods based on communicating second-order statistics. We show how these estimated class covariances can be used to initialize a linear classifier, thus exploiting the covariances without actually sharing them. When compared to state-of-the-art methods which also share only class means, our approach improves performance in the range of 4-26% with exactly the same communication cost. Moreover, our method achieves performance competitive or superior to sharing second-order statistics with dramatically less communication overhead. Finally, using our method to initialize classifiers and then performing federated fine-tuning yields better and faster convergence. Code is available at https://github.com/dipamgoswami/FedCOF.
@@ 1 Introduction
Federated learning (FL) is a widely used paradigm for distributed learning from multiple clients or participants. In FL, each client trains their local model on their private data and then send model updates to a common global server that aggregates this information into a global model. The objective is to learn a global model that performs similarly to a model jointly trained on all the client data. A major concern in existing federated optimization algorithms (McMahan et al., 2017) is the poor performance when the client data is not identically and independently distributed (iid) or when classes are imbalanced between clients (Zhao et al., 2018; Li et al., 2019; Acar et al., 2021; Karimireddy et al., 2020a). Luo et al. (2021) showed that client drift in FL is mainly due to the drift in client classifiers which optimize to the local data distribution, resulting in forgetting knowledge from other clients from previous rounds (Legate et al., 2023b; Caldarola et al., 2022). Another challenge in FL is the partial participation of clients in successive rounds (Li et al., 2019), which becomes particularly acute with large numbers of clients (Ruan et al., 2021; Kairouz et al., 2021). To address these challenges, recent works have focused on algorithms to better tackle data heterogeneity between clients  (Luo et al., 2021; Tan et al., 2022b; Legate et al., 2023a; Fanì et al., 2024).
@@
Motivated by results from transfer learning (He et al., 2019), several recent works on FL have studied the impact of using pre-trained models and observe that it can significantly reduce the impact of data heterogeneity (Legate et al., 2023a; Nguyen et al., 2023; Tan et al., 2022b; Chen et al., 2022; Qu et al., 2022; Shysheya et al., 2022; Luo et al., 2021; Tan et al., 2022a). An important finding in several of these works is that sending local class means to the server instead of raw features is more efficient in terms of communication costs, eliminates privacy concerns, and is robust to gradient-based attacks (Chen et al., 2022; Zhu et al., 2019). Tan et al. (2022b) used pre-trained models to compute and then share class means as the representative of each class, and Legate et al. (2023a) showed that aggregating local means into global means and setting them as classifier weights (FedNCM) achieves very good performance without any training. FedNCM incurs very little communication cost and enables stable initialization. Recently, the authors of Fed3R (Fanì et al., 2024) explored the impact of sharing second-order feature statistics from clients to server to solve the ridge regression problem (Boyd & Vandenberghe, 2004) in federated learning and improves over FedNCM.
Figure 1: Performance vs. communication cost using pre-trained MobileNetv2 on iNaturalist-Users-120K. Our method (FedCOF) achieves better accuracy than Fed3R while having the same cost as FedNCM.
@@
Fed3R communicates second-order statistics computed from local features for classifier initialization, and Luo et al. (2021) previously proposed using class means and covariances from all clients for classifier calibration after federated optimization. Although it is evident that exploiting second-order feature statistics results in better and more stable classifiers, it poses new problems. Notably, transferring second-order statistics for high-dimensional features from clients to the server significantly increases the communication overhead and also exposes clients to privacy risks (Luo et al., 2021; Fanì et al., 2024). In order to reap the benefits of second-order client statistics, while at the same time mitigating these risks, we propose Federated learning with COvariances for Free (FedCOF) which only communicates class means from clients to the server. We show that, from just these class means and exploiting the mathematical relationship between the covariance of class means and the class covariance matrices, we can compute an unbiased estimator of global class covariances on the server. Finally, we set the classifier weights in terms of aggregated class means and our estimated class covariances.
In this paper we exploit pre-trained feature extractors and propose a training-free method (FedCOF) that uses the same communication budget as FedNCM while delivering performance comparable to or even superior to Fed3R. FedCOF is based on a provably unbiased estimator of class covariances that requires only class means communicated from clients to the server. We validate our proposed method across several FL benchmarks, including the real-world non-iid iNaturalist-Users-120K, and our results (see Fig. 1) demonstrate that – with only a fraction of the communication costs incurred by methods based on communication of second-order statistics – FedCOF can achieve state-of-the-art results. Furthermore, we show that FedCOF can be used as an initialization for federated optimization methods in order to achieve faster and better convergence.
@@ 2 Related Work
Federated learning. FL is a rapidly-growing field and focuses on neural network training in federated environments (Zhang et al., 2021; Wen et al., 2023). Initial works like FedAvg (McMahan et al., 2017) focuses on the aspect of distributed model training and communication. Later works focus more on non-iid settings, where data among the clients is more heterogeneous (Li et al., 2019; Kairouz et al., 2021; Wang et al., 2021; Li et al., 2021). FedNova (Wang et al., 2020) normalizes local updates before averaging to address objective inconsistency. Scaffold (Karimireddy et al., 2020b) employs control variates to correct drift in local updates. FedProx (Li et al., 2020) introduces a proximal term in local objectives to stabilize the learning process. Reddi et al. (2020) proposed use of adaptive optimization methods, such as Adagrad, Adam and Yogi, at the server side. While CCVR (Luo et al., 2021) proposed a classifier calibration by aggregating class means and covariances from clients, (Li et al., 2023; Dong et al., 2022; Oh et al., 2021; Kim et al., 2024) proposed using a fixed classifier motivated by the neural collapse phenomenon (Papyan et al., 2020). After federated training with fixed classifiers, FedBABU (Oh et al., 2021) proposed to finetune the classifiers and SphereFed (Dong et al., 2022) proposed a closed-form classifier calibration.
FL with pretrained models. Recently, there has been increasing interest in incorporating pre-trained, foundation models into federated learning. Multiple works (Nguyen et al., 2023; Tan et al., 2022b; Chen et al., 2022; Qu et al., 2022; Shysheya et al., 2022) propose using pre-trained weights which reduces the impact of client data heterogeneity and achieves faster model convergence. FedFN (Kim et al., 2023) recently combined feature normalization with FedAvg to enhance performance and highlighted that using pre-trained weights can sometimes negatively affect in arbitrary heterogeneous settings. FedNCM (Legate et al., 2023a) share clients’ class means for global classifier initialization without any training. Recently, Fed3R (Fanì et al., 2024) to initialize the global classifier by using second-order feature statistics from all clients. Our work use pre-trained models and propose how we can estimate class covariances from only client means and exploit them.
@@ 3 Preliminaries
Here, we introduce the general FL problem and the training-free approach via classifier initialization.
3.1 Problem Formulation
In the FL setting, we assume $K$ clients have local datasets $D_{k}=(X_{k},Y_{k})$, where $k\in\{1,...,K\}$. We denote the total number of images from all clients as $N$ where $N=\sum_{k=1}^{K}M_{k}$ and $M_{k}$ refers to the number of images in client $k$. We represent the model as $h_{W}(f_{\theta}(x))$ which can be decomposed into two parts: the feature extractor $f$ parameterized by $\theta$ which gives a $d$-dimensional embedding from a given image and the final classifier layer $h:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$ parameterized by $W$ where $C$ refers to the total number of classes. The general objective of federated optimization (Konečnỳ et al., 2016) is to learn a global model that minimizes the sum of the losses across all the clients as follows:
$\operatorname*{arg\,min}\limits_{\theta,W}\sum_{k=1}^{K}\frac{M_{k}}{N}\mathcal{L}(h_{W}(f_{\theta}(X_{k})),Y_{k})$ (1)
where $\mathcal{L}$ is the classification loss function (e.g., cross-entropy). With the growing quality of pretrained models, recent works on FL (Chen et al., 2022; Legate et al., 2023a; Nguyen et al., 2023; Tan et al., 2022b; Fanì et al., 2024) has focused on scenarios where all clients start with a strong pre-trained network. After initializing $\theta$ with pre-trained weights, the models can be optimized in federated manner. In this work, we propose to use the frozen pre-trained model without performing local updates across clients.
3.2 Training-free Federated Learning Methods
Federated NCM. Legate et al. (2023a) propose a Nearest Class Mean (NCM) classifier where the global classifier weights (considering a linear layer) for class $c$ denoted by $W_{c}$ can be initialized with normalized global class means $\hat{\mu}_{c}$, which are aggregated from the local class means $\hat{\mu}_{k,c}$ as follows:
$W_{c}=\frac{\hat{\mu}_{c}}{\|\hat{\mu}_{c}\|};\quad\quad\hat{\mu}_{c}=\frac{1}{N_{c}}\sum_{k=1}^{K}n_{k,c}\,\hat{\mu}_{k,c};\quad\quad\hat{\mu}_{k,c}=\frac{1}{n_{k,c}}\sum_{x\in X_{k,c}}f(x)$ (2)
where $X_{k,c}$ is the subset of $X_{k}$ containing only images of class $c$ , $n_{k,c}$ refers to number of images in $X_{k,c}$ and $N_{c}=\sum_{k=1}^{K}n_{k,c}$ is the total number of images of class $c$ across all clients. FedNCM needs to communicate class means $\hat{\mu}_{k,c}$ and class counts $n_{k,c}$ from clients to the server only once.
Federated Ridge Regression. While FedNCM exploits only class means, Fed3R (Fanì et al., 2024) recently proposed to use ridge regression which needs second-order feature statistics from all clients to initialize the global classifier, leading to improved performance compared to FedNCM. The ridge regression problem aims to find the optimal weights that minimize the following objective:
@@ 3.2 Training-free Federated Learning Methods
where $F\in\mathbb{R}^{d\times N}$ is the feature matrix extracted from a pre-trained model and $Y\in\mathbb{R}^{N\times C}$ contains one-hot encoding labels for the $N$ features with $C$ classes. The closed-form solution is given by:
@@
with $G=FF^{\top}\in\mathbb{R}^{d\times d}$ and $B=FY\in\mathbb{R}^{d\times C}$, $\lambda\in\mathbb{R}$ is an hyper-parameter and $I_{d}$ is the $d\times d$ identity matrix.
In Fed3R (Fanì et al., 2024), each client $k$ computes two local matrices $G_{k}=F_{k}F_{k}^{\top}\in\mathbb{R}^{d\times d}$ and $B_{k}=F_{k}Y_{k}\in\mathbb{R}^{d\times C}$, where $F_{k}$ and $Y_{k}$ are the feature matrix and the labels of client $k$, and then sends them to the global server. The server aggregates these matrices as
$G=\sum_{k=1}^{K}G_{k},\quad B=\sum_{k=1}^{K}B_{k}$ (5)
and compute $W^{*}$ (Eq.4), which is normalized and then used to initialize the global linear classifier.
Table 1: FedNCM (Legate et al., 2023a) shares only class means $\hat{\mu}_{k,c}$ and has minimal communication. Fed3R (Fanì et al., 2024) requires sum of class features $B_{k}$ and feature matrix $G_{k}$ from all clients, thereby increasing the communication budget by $d^{2}K$. We propose FedCOF, which shares only class means and estimates a global class covariance $\hat{\Sigma}_{c}$ to initialize the classifier weights.
Classifier Initialization Each Client Shares Server Uses Communication Cost
FedNCM (Legate et al., 2023a) $\{\hat{\mu}_{k,c}\}_{c=1}^{C}$$\{\{\hat{\mu}_{k,c}\}_{c=1}^{C}\}_{k=1}^{K}$ $dCK$
Fed3R (Fanì et al., 2024) $G_{k},B_{k}$$\{G_{k},B_{k}\}_{k=1}^{K}$ $(dC+d^{2})K$
FedCOF (ours) $\{\hat{\mu}_{k,c}\}_{c=1}^{C}$$\{\{\hat{\mu}_{k,c}\}_{c=1}^{C}\}_{k=1}^{K},\{\hat{\Sigma}_{c}\}_{c=1}^{C}$ $dCK$
@@ 4 Federated Learning with COvariances for Free (FedCOF)
Here, we derive our FedCOF method which is based on a provably unbiased estimator of class covariances and requires only class means communicated from clients to server.

### v3

# Covariances for Free: Exploiting Mean Distributions for Training-free Federated Learning

@@ Abstract
Using pre-trained models has been found to reduce the effect of data heterogeneity and speed up federated learning algorithms. Recent works have explored training-free methods using first- and second-order statistics to aggregate local client data distributions at the server and achieve high performance without any training. In this work, we propose a training-free method based on an unbiased estimator of class covariance matrices which only uses first-order statistics in the form of class means communicated by clients to the server. We show how these estimated class covariances can be used to initialize the global classifier, thus exploiting the covariances without actually sharing them. We also show that using only within-class covariances results in a better classifier initialization. Our approach improves performance in the range of 4-26% with exactly the same communication cost when compared to methods sharing only class means and achieves performance competitive or superior to methods sharing second-order statistics with dramatically less communication overhead. The proposed method is much more communication-efficient than federated prompt-tuning methods and still outperforms them. Finally, using our method to initialize classifiers and then performing federated fine-tuning or linear probing again yields better performance. Code is available at https://github.com/dipamgoswami/FedCOF.
@@ 1 Introduction
Federated learning (FL) is a widely used paradigm for distributed learning across multiple clients. In FL, each client trains their local model on their private data and then sends model updates to a common global server that aggregates this information into a global model. The objective is to learn a global model that performs similarly to a model jointly trained on all the client data. A major concern in existing FL algorithms (mcmahan2017communication, ) is the poor performance when the client data is not identically and independently distributed (iid) or when classes are imbalanced between clients (zhao2018federated, ; li2019convergence, ; acar2021federated, ; karimireddy2020scaffold, ). In luo2021no , the authors showed that client drift in FL is mainly due to drift in client classifiers which optimize to local data distributions, resulting in forgetting knowledge from clients of previous rounds (legate2023re, ; caldarola2022improving, ). Another challenge in FL is the partial participation of clients in successive rounds (li2019convergence, ), which becomes particularly acute with large numbers of clients (ruan2021towards, ; kairouz2021advances, ). To address these challenges, recent works have focused on algorithms to better tackle data heterogeneity across clients (luo2021no, ; tan2022federated, ; legate2023guiding, ; fani2024accelerating, ).
@@
Motivated by results from transfer learning (He_2019_ICCV, ), several recent works on FL have studied the impact of using pre-trained models and observe that it can significantly reduce the impact of data heterogeneity (legate2023guiding, ; nguyen2022begin, ; tan2022federated, ; chen2022importance, ; qu2022rethinking, ; shysheya2022fit, ; luo2021no, ; tan2022fedproto, ; weng2024probabilistic, ). An important finding in several of these works is that sending local class means to the server instead of raw features is more efficient in terms of communication costs, eliminates privacy concerns, and is robust to gradient-based attacks (chen2022importance, ; zhu2019deep, ). The authors of tan2022federated used pre-trained models to compute and then share class means as the representative of each class, and in legate2023guiding the authors showed that aggregating local means into global means and setting them as classifier weights (FedNCM) achieves very good performance without any training. FedNCM incurs very little communication cost and enables stable initialization. Recently, the authors of Fed3R (fani2024accelerating, ) explored the impact of sharing second-order feature statistics from clients to server to solve the ridge regression problem (boyd2004convex, ) in FL and improve over FedNCM. The sharing of second-order statistics has also previously been explored for classifier calibration after federated optimization (luo2021no, ).
@@
Although it is evident that exploiting second-order feature statistics results in better and more stable classifiers, it poses new problems. Notably, sharing second-order statistics for high-dimensional features from clients to the server significantly increases communication overhead and exposes clients to privacy risks (luo2021no, ; fani2024accelerating, ). To reap the benefits of second-order client statistics, while at the same time mitigating these risks, in this paper we propose Federated learning with COvariances for Free (FedCOF) which only communicates class means from clients to the server (see Section˜1). We show that, from just these class means and the mathematical relationship between their covariance and the class covariance matrices, we can compute a provably unbiased estimator of global class covariances on the server. FedCOF is a training-free method that exploits pre-trained feature extractors. It uses the same communication budget as FedNCM while delivering performance comparable to or even superior to Fed3R (see Figure˜1). Additionally, we show how to improve classifier initialization using only within-class covariances, setting the classifier weights based on aggregated class means and our estimated class covariances. To summarize, our main contributions are:
  • We propose FedCOF, a training-free method that exploits a provably unbiased estimator of class covariances, which requires only class means from clients, thus avoiding the need to share second-order statistics, reducing communication costs and mitigating privacy concerns.
  • We show how FedCOF leverages within-class covariances, estimated solely from client means, to initialize the global classifier, yielding more stable and better-conditioned solutions than using both within- and between-class scatter matrices.
  • We validate FedCOF across multiple FL benchmarks, including the non-iid iNaturalist-Users-120K dataset, achieving state-of-the-art results with lower communication cost than methods using second-order statistics, and showing superior performance to recent federated prompt-tuning approaches while also serving as an effective initialization for subsequent federated optimization methods such as fine-tuning and linear probing.
Table 1: FedNCM (legate2023guiding, ) shares only class means $\hat{\mu}_{k,c}$ and has minimal communication. Fed3R (fani2024accelerating, ) requires sum of class features $B_{k}$ and feature matrix $G_{k}$ from all clients, thereby increasing the communication cost by $d^{2}K$. We propose FedCOF, which shares only class means and estimates a global class covariance $\hat{\Sigma}_{c}$ to initialize the classifier weights. Note that only a small subset of all classes are present in each client. For simplicity, we show the upper bound of communication cost here where $C^{\prime}$ denotes the maximum number of classes present in a single client.
Method Client Shares Server Uses Comm. Cost
FedNCM $\{\hat{\mu}_{k,c},n_{k,c}\}_{c=1}^{C^{\prime}}$ $\{\{\hat{\mu}_{k,c},n_{k,c}\}_{c=1}^{C^{\prime}}\}_{k=1}^{K}$ $dC^{\prime}K$
Fed3R $G_{k},B_{k}$ $\{G_{k},B_{k}\}_{k=1}^{K}$ $(dC^{\prime}+d^{2})K$
FedCOF $\{\hat{\mu}_{k,c},n_{k,c}\}_{c=1}^{C^{\prime}}$ $\{\{\hat{\mu}_{k,c},n_{k,c}\}_{c=1}^{C^{\prime}}\}_{k=1}^{K},\{\hat{\Sigma}_{c}\}_{c=1}^{C^{\prime}}$ $dC^{\prime}K$
Figure 1: Accuracy vs. communication cost using pre-trained MobileNetv2 on iNaturalist-Users-120K. FedCOF outperforms Fed3R while having the same communication cost as FedNCM.
@@ 2 Related Work
Federated learning. FL focuses on neural network training in distributed environments (zhang2021survey, ; wen2023survey, ). Initial works like FedAvg (mcmahan2017communication, ) proposed training by averaging of distributed models. Later works focus more on non-iid settings, where data among the clients is more heterogeneous (li2019convergence, ; kairouz2021advances, ; wang2021field, ; lifedbn, ). FedNova (Wang2020fednova, ) normalizes local updates before averaging to address objective inconsistency. Scaffold (karimireddy20scaffold, ) employs control variates to correct drift in local updates. FedProx (li2020fedprox, ) introduces a proximal term in local objectives to stabilize the learning process. reddi2020adaptive proposed use of adaptive optimization methods, such as Adagrad, Adam and Yogi, at the server side. While CCVR (luo2021no, ) proposed a classifier calibration method that aggregates class means and covariances from clients, other recent works li2023no ; dong2022spherefed ; oh2021fedbabu ; kim2024feddr+ proposed using fixed classifiers inspired by the neural collapse phenomenon. After federated training with fixed classifiers, FedBABU (oh2021fedbabu, ) proposed to fine-tune the classifiers and SphereFed (dong2022spherefed, ) proposed a closed-form classifier calibration.
FL with pre-trained models. While conventional FL methods start training from scratch without any pre-training, we focus on the FL setting using pre-trained models. FedFN (kim2023fedfn, ) recently highlighted that using pre-trained weights can sometimes negatively impact performance. However, there has been increasing interest in incorporating pre-trained, foundation models into federated learning. Multiple works propose using pre-trained weights which reduces the impact of client data heterogeneity and achieves faster model convergence (nguyen2022begin, ; tan2022federated, ; chen2022importance, ; qu2022rethinking, ; shysheya2022fit, ; weng2024probabilistic, ). Federated prompt-tuning methods (weng2024probabilistic, ) proposed to reduce the communication cost by tuning only prompts in a federated manner using pre-trained ViT models. Recently, it has been shown that training-free methods using pre-trained networks, achieve strong performance without any training by exploiting feature class means (legate2023guiding, ) or second-order feature statistics (fani2024accelerating, ). Another recent work beitollahifoundation proposed sharing Gaussian mixture models (multiple sets of means and covariances) for each class from clients to server for one-shot FL. In this work, we propose a training-free method with pre-trained models that estimates class covariances from only client means for initializing the global classifier.
@@ 3 Preliminaries
In the FL setting we assume $K$ clients, each with a local dataset $D_{k}=(X_{k},Y_{k})$, for $k\in\{1,...,K\}$, and denote by $N$ the total number of images across all clients. The model is composed of a feature extractor $f$, parameterized by $\theta$, which maps images to $d$-dimensional embeddings, and a classifier $h:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$, parameterized by $W$, where $C$ is the total number of classes. In federated optimization, each client trains its local model on its private data $D_{k}$ and transmits only intermediate information – such as parameter updates or feature statistics – while a central server aggregates these signals to minimize a global objective, without revealing raw data (konevcny2016federated, ).
With the growing availability of high-quality pre-trained models, recent works have focused on scenarios in which all clients are initialized with the same pre-trained feature extractor  (chen2022importance, ; legate2023guiding, ; nguyen2022begin, ; tan2022federated, ; fani2024accelerating, ; weng2024probabilistic, ). Notably, training-free methods (fani2024accelerating, ; legate2023guiding, ) – in which only the global classifier is initialized using client feature statistics, without training the feature extractor – achieve strong performance. They often outperform federated full fine-tuning methods, such as FedAdam and FedAvg, which train client backbones and classifiers, starting from the same shared pre-trained feature extractor but with randomly initialized classifiers, at a fraction of the communication and computation costs. Moreover, this training-free initialization can serve as an effective starting point, improving the performance of federated full fine-tuning. We now discuss recent training-free methods.
Federated NCM. FedNCM legate2023guiding employs a Nearest Class Mean (NCM) classifier, where the global linear classifier weights for class $c$ (denoted as $W_{c}$) are initialized as $\hat{\mu}_{c}/\|\hat{\mu}_{c}\|$. Here, $\hat{\mu}_{c}$ represents the global class mean aggregated from the local client class means $\hat{\mu}_{k,c}$ as follows:
$\hat{\mu}_{c}=\frac{1}{N_{c}}\sum_{k=1}^{K}n_{k,c}\,\hat{\mu}_{k,c};\quad\hat{\mu}_{k,c}=\frac{1}{n_{k,c}}\sum_{x\in X_{k,c}}f(x),$ (1)
where $X_{k,c}$ is a subset of $X_{k}$ having images of class $c$, $n_{k,c}$ refers to number of images in $X_{k,c}$, and $N_{c}=\sum_{k=1}^{K}n_{k,c}$ is the number of images of class $c$ across all clients.
Federated Ridge Regression. While FedNCM exploits only class means, Fed3R (fani2024accelerating, ) recently proposed using ridge regression which requires second-order feature statistics from all clients to initialize the global classifier, leading to improved performance compared to FedNCM. The ridge regression problem aims to find the optimal weights that minimize the following objective:
@@
where $F\in\mathbb{R}^{d\times N}$ is the feature matrix extracted using a pre-trained model and $Y\in\mathbb{R}^{N\times C}$ contains one-hot encoded labels for the $N$ features over $C$ classes. The closed-form solution is given by:
@@
where $G=FF^{\top}\in\mathbb{R}^{d\times d}$, $B=FY\in\mathbb{R}^{d\times C}$, $I_{d}\in\mathbb{R}^{d\times d}$ is the identity matrix, and $\lambda\in\mathbb{R}$ is a hyperparameter. In Fed3R, each client $k$ computes two local matrices $G_{k}=F_{k}F_{k}^{\top}\in\mathbb{R}^{d\times d}$ and $B_{k}=F_{k}Y_{k}\in\mathbb{R}^{d\times C}$, where $F_{k}$ and $Y_{k}$ are the feature matrix and the labels of client $k$, and then sends them to the global server. The server aggregates these matrices as $G=\sum_{k=1}^{K}G_{k},\quad B=\sum_{k=1}^{K}B_{k}$ to compute $W^{*}$, which is normalized and used to initialize the global classifier.
@@ 4 Federated Learning with COvariances for Free (FedCOF)
In this section, we derive our FedCOF method which is based on a provably unbiased estimator of class covariances and requires transfer of only class means from clients to server.